# Energy Optimized Piecewise Polynomial Approximation Utilizing Modern Machine Learning Optimizers*

### v1

@@ Abstract
This work explores an extension of Machine Learning (ML)-optimized Piecewise Polynomial (PP) approximation by incorporating energy optimization as an additional objective. Traditional closed-form solutions enable continuity and approximation targets but lack flexibility in accommodating complex optimization goals. By leveraging modern gradient descent optimizers within TensorFlow, we introduce a framework that minimizes total curvature in cam profiles, leading to smoother motion and reduced energy consumption for input data that is unfavorable for sole approximation and continuity optimization. Experimental results confirm the effectiveness of this approach, demonstrating its potential to improve efficiency in scenarios where input data is noisy or suboptimal for conventional methods.
Index Terms:
@@
Piecewise Polynomials, Gradient Descent Optimizers, Approximation, TensorFlow, Electronic Cams, Energy Consumption
I INTRODUCTION
I-A Motivation and Problem Statement
@@
PPs are commonly utilized in various scientific and engineering disciplines. One key area of interest is trajectory planning for machines in mechatronics, computing time-dependent position, velocity, or acceleration profiles that provide setpoints for controllers of industrial machines. In this context, electronic cams define the repetitive motion of servo drives, commonly defined as input point clouds and approximated by PPs. Traditionally, the approximation process involves solving a system of equations in closed form, incorporating domain-specific constraints like continuity, cyclicity, or periodicity. While computationally efficient, closed-form solutions offer limited flexibility regarding polynomial degrees, bases, or the integration of additional constraints.
@@
As optimization goals grow more complex, numerical methods such as gradient descent become beneficial. Since gradient-based optimization lies at the heart of training neural networks, modern ML frameworks like TensorFlow or PyTorch come with a range of state-of-the art optimizers. In previous work [5] [1], we utilize these optimizers directly to fit $C^{k}$-continuous PPs for the use in the electronic cam domain. The problem formulation allows for the addition of further domain-specific optimization targets, however, it is not clear whether results converge without introducing further regularization measures.
This is why, in this work, we investigate optimizing the total curvature of PP cam curves as an additional domain-specific goal. This holds the potential of leading to smoother motion profiles and therefore improved efficiency in machines in cases where the input data is noisy or unfavorable for continuity and approximation optimization alone, leading to reduced forces and energy consumption in such scenarios.
I-B Contribution
Considering the $3$-objective optimization target described in section II-B, the set of Pareto efficient solutions for input data unfavorable for sole continuity and approximation optimization can be sketched as depicted in fig. 1. This concerns input data where results end up in the left upper corner of feasible solutions. In this work, we demonstrate that an extension of the framework introduced in [5] and [1] towards energy optimization is feasible and indeed improves results in the case of such input data.
Figure 1: Sketch of Pareto frontier (dashed) for this $3$-objective optimization. Feasible solutions form the cross-hatched area.
I-C Prior and Related Work
@@
II PROBLEM FORMULATION
II-A Model Definition
Input data consists of $n$ samples $x_{1}\leq\dots\leq x_{n}\in\mathbb{R}$ with $y_{i}\in\mathbb{R}$. We define the polynomial boundaries of a PP function $f$ on the interval $I=[x_{1},x_{n}]$ as $\xi_{0}\leq\dots\leq\xi_{m}$, with $\xi_{0}=x_{1}$ and $\xi_{m}=x_{n}$. Each of the $m$ segments $p_{i}\colon I\to\mathbb{R}$ is modeled via polynomials of degree $d$ as
$p_{i}=\sum_{j=0}^{d}\alpha_{i,j}x^{j}.$ (1)
The coefficients $\alpha_{i,j}$ are the model parameters trained via TensorFlow gradient descent optimizers according to the loss function described in the next section.
II-B Loss Function
In order to reduce the forces applied to the cam-follower system, we want to optimize the cam’s acceleration curve. To do so, we can quantify total curvature as $\int_{I}f^{\prime\prime}(x)^{2}\;\mathrm{d}{x}\,$ for a cam profile $f$ over the interval $I$. Considering our individual polynomial segments $p_{i}$, we receive
$\int_{I}f^{\prime\prime}(x)^{2}\;\mathrm{d}{x}\,$ $=\sum_{i=1}^{m}\int_{I_{i}}p^{\prime\prime}_{i}(x)^{2}\;\mathrm{d}{x}\,$ (2)
with
@@ II-B Loss Function
Plugging in the boundary points for the definite integral and generalizing to the overall PP with $m$ segments then leaves us with the energy loss
$\ell_{E}=\sum_{i=1}^{m}\sum_{j=2}^{d}\sum_{k=2}^{d}\alpha_{i,j}\alpha_{i,k}\frac{jk(j-1)(k-1)}{j+k-3}\cdot\left(\xi_{i}^{j+k-3}-\xi_{i-1}^{j+k-3}\right).$ (5)
We quantify the approximation error $\ell_{2}$ as
$\ell_{2}=\frac{1}{n}\sum_{i}|f(x_{i})-y_{i}|^{2}$ (6)
and sum up discontinuities at all $\xi_{i}$ across relevant derivatives for $C^{k}$-continuity as
$\ell_{\text{CK}}=\frac{1}{m-1}\sum_{i=1}^{m-1}\sum_{j=0}^{k}\left(\frac{\Delta_{i,j}}{r_{k}}\right)^{2}\text{with}\Delta_{i,j}=p^{(j)}_{i+1}(\xi_{i})-p^{(j)}_{i}(\xi_{i})$ (7)
and achieve cyclicity by adapting $m-1$ to $m$ and generalizing $\delta_{i,j}=p^{(j)}_{1+(i\bmod m)}(\xi_{i\bmod m})-p^{(j)}_{i}(\xi_{i})$, ignoring the case $j=0$ when $i=m$. We then can prioritize either of the three objectives by defining the weighted total loss as
$\ell=\alpha\ell_{\text{CK}}+\beta\ell_{2}+(1-\alpha-\beta)\ell_{\text{E}},$ (8)
where $\alpha,\beta\geq 0$ and $\alpha+\beta\leq 1$.
III EXPERIMENTAL RESULTS
In an effort of shifting non-$\ell_{\text{E}}$-optimized results to the upper left corner in the sense of fig. 1, we define $100$ data points $x\in[0,1]$ and let $y=\sin{(x^{2}4\pi)}$. We then add noise to the data via NumPy’s random.normal function with a scale of $0.1$. In practice, such data could stem from measurements, introducing some noise to the input cam data. We utilize TensorFlow optimizers to fit our introduced PP model to this data. We run $1000$ epochs with early stopping and patience of $100$ epochs.
To have enough free parameters, we fit a $C^{2}$-continuous PP of degree $7$. In addition to the nature of the data, in an effort of further shifting results to the upper left corner in the sense of fig. 1, each segment should not have too many data points, i.e. the number of polynomial segments needs to be sufficiently high. Table I shows the result of fitting a $16$ segment PP with and without $\ell_{\mathrm{E}}$ optimization utilizing the AMSGrad version of Adam in this context.
TABLE I: Results for fitting a $C^{2}$-continuous PP of degree $7$.
Parameters $\ell_{2}$ $\ell_{\mathrm{CK}}$ $\ell_{\mathrm{E}}$
$\alpha=0.9,\beta=0.1$ $0.0079$ $0.0026$ $5.3551$
$\alpha=0.9,\beta=0.099$ $0.0096$ $0.0210$ $0.9326$
Table I confirms that optimizing $\ell_{\text{E}}$ in addition to $\ell_{2}$ and $\ell_{\text{CK}}$ works. Oscillations in the curve’s shape are reduced, therefore providing a smoother result. This can be observed from plots of the resulting PPs along with further experiments available at [4].
IV SUMMARY AND OUTLOOK
@@
Our results demonstrate that an extension of the framework introduced in [5] and [1] towards energy optimization is feasible. Results converge without additional regularization measures and indeed improve in the case of input data that is unfavorable for sole approximation and continuity optimization.
Further work could focus on hyperparameter optimization or measures to improve performance of $\ell_{\text{E}}$ optimization in combination with orthogonal bases.

### v3

@@ Abstract
This work explores an extension of Machine Learning (ML)-optimized Piecewise Polynomial (PP) approximation by incorporating energy optimization as an additional objective. Traditional closed-form solutions enable continuity and approximation targets but lack flexibility in accommodating complex optimization goals. By leveraging modern gradient descent optimizers within TensorFlow, we introduce a framework that minimizes elastic strain energy in cam profiles, leading to smoother motion. Experimental results confirm the effectiveness of this approach, demonstrating its potential to Pareto-efficiently trade approximation quality against energy consumption.
Keywords:
@@
Piecewise Polynomials Gradient Descent Approximation TensorFlow Electronic Cams Energy Optimization
1 Introduction
1.1 Motivation and Problem Statement
@@
PPs are commonly utilized in various scientific and engineering disciplines. One key area of interest is trajectory planning for machines in mechatronics, computing time-dependent position, velocity, or acceleration profiles that provide setpoints for controllers of industrial machines. In this context, electronic cams define the repetitive motion of servo drives, commonly defined as input point clouds and approximated by PPs. Traditionally, the approximation process involves solving a system of equations in closed form, incorporating domain-specific constraints like continuity, cyclicity, or periodicity. While computationally efficient, closed-form solutions offer limited flexibility regarding polynomial degrees, polynomial bases, or the integration of additional constraints.
@@
As optimization goals grow more complex, numerical methods such as gradient descent become beneficial. Since gradient-based optimization lies at the heart of training (deep) neural networks, modern ML frameworks like TensorFlow or PyTorch come with a range of recent optimizers. In previous work [5] [1], we utilize these optimizers directly to fit $\mathcal{C}^{k}$-continuous PPs for the use in the electronic cam domain.
In this work we extend the problem setting to a constrained multi-objective optimization problem by adding an energy term as a domain-specific loss term and investigate (i) convergence and (ii) the pareto front under this extension. More precisely, given input points $x_{1}\leq\dots\leq x_{n}\in\mathbb{R}$ and respective target values $y_{1},\dots,y_{n}\in\mathbb{R}$, we look for a PP function $f$ that minimizes both the approximation loss $\ell_{2}$ and energy loss $\ell_{E}=\int_{I}f^{\prime\prime}(x)^{2}\mathrm{d}{x}\,$ over some domain $I$ that captures the total curvature, under the constraint that $f\in\mathcal{C}^{k}$, where $\mathcal{C}^{k}$ denotes the set of $k$-times continuously differentiable functions $\mathbb{R}\to\mathbb{R}$. We form a loss $\ell_{\text{CK}}$ that captures the “amount of violation” of $\mathcal{C}^{k}$-continuity, which will lead to the optimization constraint $\ell_{\text{CK}}=0$, as it is of practical relevance to feed servo drives with PPs of continuity class $\mathcal{C}^{k}$ to prevent excessive forces, vibrations and wear on machine parts.
Note that $\ell_{E}$ is related to energy in multiple ways: In a mechanical setting, $\ell_{E}$ is the elastic strain energy of a rod bent as the function graph of $f$. When $f$ is used as a cam and we approximately assume that the acceleration $f^{\prime\prime}$ is proportional to the motor coil current then $\ell_{E}$ is proportional to the copper losses in the motor coils.
1.2 Prior and Related Work
@@
2 The optimization problem
Parameterized model.
Let us consider $n$ input points $x_{1}\leq\dots\leq x_{n}\in\mathbb{R}$ with respective target values $y_{1},\dots,y_{n}\in\mathbb{R}$. We split the input domain $I=[x_{1},x_{n}]$ into $m$ sub-intervals $I_{i}=[\xi_{i-1},\xi_{i}]$, with $1\leq i\leq m$, where $x_{1}=\xi_{0}\leq\dots\leq\xi_{m}=x_{n}$. A PP function $f$ of degree $d$ is defined by $m$ polynomial functions $p_{1},\dots,p_{m}$ of degree $d$ as
$p_{i}(x)=\sum_{j=0}^{d}\alpha_{i,j}x^{j},$ (1)
where $p_{i}$ covers the sub-domain $I_{i}$ of $I$. As a parameterized model, $f$ possesses $(d+1)\cdot m$ model parameters $\alpha_{i,j}$, which are trained via gradient descent optimizers according to the loss functions described below.
Loss functions.
We build on [1] for the approximation loss $\ell_{2}$ and the $\mathcal{C}^{k}$-continuity loss $\ell_{\text{CK}}$. The former is defined as
$\ell_{2}=\frac{1}{n}\sum_{i=1}^{n}|f(x_{i})-y_{i}|^{2}$ (2)
and the latter introduces a term $\Delta_{i,j}$ that measures the discontinuity of the $j$-th derivative at $\xi_{i}$ to then define
$\ell_{\text{CK}}=\frac{1}{m-1}\sum_{i=1}^{m-1}\sum_{j=0}^{k}\Delta_{i,j}^{2}\quad\text{with}\quad\Delta_{i,j}=p^{(j)}_{i+1}(\xi_{i})-p^{(j)}_{i}(\xi_{i}).$ (3)
Recall that the $\xi_{i}$ are the boundary points of the domains $I_{i}$ of $p_{i}$. Note that achieving $\ell_{\text{CK}}=0$ means that $f$ is $\mathcal{C}^{k}$-continuous and $\ell_{2}=0$ means zero approximation error.
We can account for cyclicity (e.g., of the cam profile) by replacing in eq.˜3 the two occurrences of m1m-1 by $m$ and generalizing $\Delta_{i,j}$ to
$\Delta_{i,j}=p^{(j)}_{1+(i\bmod m)}(\xi_{i\bmod m})-p^{(j)}_{i}(\xi_{i}).$ (4)
(We may exclude the case $j=0$ when $i=m$ if we do not require periodicity, see [1] for details.)
Energy loss.
In this work, we motivated the energy loss $\ell_{E}$ for $f$ as
$\ell_{\mathrm{E}}=\int_{I}f^{\prime\prime}(x)^{2}\;\mathrm{d}{x}\,$ $=\sum_{i=1}^{m}\int_{I_{i}}p^{\prime\prime}_{i}(x)^{2}\;\mathrm{d}{x}\,$ (5)
in the introduction and further note that
@@ 2 The optimization problem
Solving the definitive integral in eq.˜6 and plugging it into eq.˜5 gives us the energy loss in closed form as a function of the trainable model parameters:
$\ell_{\mathrm{E}}=\sum_{i=1}^{m}\sum_{j=2}^{d}\sum_{k=2}^{d}\alpha_{i,j}\alpha_{i,k}\frac{jk(j-1)(k-1)}{j+k-3}\cdot\left(\xi_{i}^{j+k-3}-\xi_{i-1}^{j+k-3}\right).$ (7)
Multi-objective optimization.
[…]
3 Experimental Results and Discussion
The code and results discussed in this section along with further experiments and plots are provided in the public repository [4].
For our experiment, we consider a dataset with $n=100$ input points $x_{i}$ with target values $y_{i}=\sin(4\pi x_{i}^{2})+n_{i}$, where $n_{i}$ is a normal noise with zero mean and standard deviation $0.1$, produced by numpy.random.normal. In practice, such data could stem from measurements. In fig.˜1 the dataset is illustrated and we observe that solely optimizing for approximation loss leads to significant curvatures (left), while adding $\ell_{\mathrm{E}}$ regularizes the result (right).
Figure 1: Results for the first two rows of table˜1. Left: $\beta=0.9$, right: $\beta=0.45$.
We utilize the AMSGrad optimizer in TensorFlow to fit our PP model as explained in section˜2. Within our training loop, we run $1000$ epochs with early stopping and patience of $100$ epochs to fit a $\mathcal{C}^{2}$-continuous, periodic PP of degree $7$ with $16$ polynomial segments. Table˜1 and fig.˜2 show the results after establishing $\ell_{\text{CK}}=0$ using algorithm CKMIN from [5] for different $\alpha$ and $\beta$. Table˜1 confirms that additionally optimizing $\ell_{\mathrm{E}}$ works: Reducing $\beta$ from $0.9$ to $0.45$, shown in the first two rows of table˜1, reduces $\ell_{\mathrm{E}}$ by a factor of $4.40$.
Figure 2: $\ell_{\mathrm{E}}$ over $\ell_{2}$ for different $\alpha$ and $\beta\in(0,1-\alpha]$. $\star$ is Pareto optimal.
Parameters $\ell_{2}$ $\ell_{\mathrm{E}}$
$\alpha=0.10,\beta=0.900$ $0.005$ $8.682$
$\alpha=0.10,\beta=0.450$ $0.014$ $1.970$
$\alpha=0.50,\beta=0.500$ $0.007$ $5.468$
$\alpha=0.50,\beta=0.250$ $0.046$ $0.616$
$\alpha=0.75,\beta=0.250$ $0.007$ $5.852$
$\alpha=0.75,\beta=0.125$ $0.077$ $0.282$
$\alpha=0.99,\beta=0.010$ $0.043$ $4.977$
$\alpha=0.99,\beta=0.005$ $0.135$ $0.121$
Table 1: Results for different $\alpha,\beta$.
4 Conclusion
@@
Our results demonstrate that an extension of the framework introduced in [5] and [1] towards energy optimization is feasible. Results converge without additional regularization measures and indeed improve in the case of input data that is unfavorable for sole approximation and continuity optimization. Optimizing $\ell_{\mathrm{E}}$ further has positive side effects: Figure˜1 shows that oscillations in the curve’s shape are reduced, which is favorable from a mechanical perspective.
While orthogonal bases, like Chebyshev polynomials, have shown much better performance on $\ell_{2}$, preliminary tests indicated convergence is impaired when adding $\ell_{\mathrm{E}}$. We leave the investigation of mitigation measures to future work.
{credits}
4.0.1 Acknowledgements
This work was supported by the Christian Doppler Research Association (JRC ISIA) and the European Interreg Österreich-Bayern project BA0100172 AI4GREEN. We want to thank the reviewers for their valuable feedback.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.